If you roll a die 100 times, what is the approximate probability that you will roll between 9 and 16 ones, inclusive? (Round your answer to two decimal places.) HINT [See Example 4.]
Answer #1

=P((8.5-16.6667)/3.7627<Z<(16.5-16.6667)/3.7627)
=P(-2.17<Z<-0.04)
=P(Z<-0.04)-P(Z<-2.17)
=0.484-0.015
=0.47